OpenStudy (katielong):

express: (4x^2+23x+24)/(x^2+5x+6) in partial fractions

OpenStudy (katielong):

\[\frac{ 4x^2+23x+24 }{ x^2+5x+6 }\]

OpenStudy (marco26):

First you need to decompose the fraction using variables like A, B, etc.

OpenStudy (katielong):

Ax + B + C/(x+2) + D/(x+3) ?

OpenStudy (katielong):

@marco26 ?

OpenStudy (marco26):

\[\frac{ 4x ^{2}+23x+24 }{ (x+2)(x+3) }=\frac{ A }{ (x+2) }+\frac{ B }{ (x+3) }\] -> this is what you need to get when you decompose the fraction.

OpenStudy (katielong):

okay, but isn't a 'C' term needed because the denominator is to the same degree as the numerator?

OpenStudy (marco26):

Sorry I was wrong. With the same degree you have to do long division first. Divide the numerator with the denominator and you should get \[\frac{ 4x ^{2} +23x+24}{ x ^{2}+5x+6 }=4+\frac{ 3x }{ x ^{2}+5x+6 }\] and proceed to decomposition.

OpenStudy (anonymous):

now it's right